OpenStudy (anonymous):

How do I resolves this: r=10 ; (h,k)=(-6,8)

OpenStudy (anonymous):

what do you want? a circle equation?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] this the center-radius form for the equation of a circle. plug in your values for h, k and r. what do you get?

OpenStudy (anonymous):

in H i put -6 and K 8 and it is = 100

OpenStudy (anonymous):

yes. you can simplify it to (x+6)^2 + (y-8)^2 = 100

OpenStudy (anonymous):

Yes, thats it..

OpenStudy (anonymous):

yw